# Another algorithm template

### v11

@@ 1 Introduction
$\|P_{n}(v)\|_{V}\leq C\|v\|_{V},\quad\|v-P_{n}(v)\|_{V}\leq Cn^{-\alpha}\|v\|_{W},\quad\forall n\in{\mathbb{N}}_{0},\quad\forall v\in W.$ (1.5)
@@ 2 Linear approximation by truncated Hermite GPC expansion
$\|P_{n}(v)\|_{X^{1}}\leq C\|v\|_{X^{1}},\quad\|v-P_{n}(v)\|_{X^{1}}\leq Cn^{-\alpha}\|v\|_{X^{2}},\quad\forall n\in{\mathbb{N}}_{0},\quad\forall v\in X^{2}.$ (2.3)
@@
$\delta_{k}(v):=\ P_{2^{k}}(v)-P_{2^{k-1}}(v),\ k\in{\mathbb{N}},\quad\delta_{0}(v)=P_{0}(v).$
We can represent every $v\in X^{2}$ by the series
@@
converging in $X^{1}$ and satisfying the estimate
$\|\delta_{k}(v)\|_{X^{1}}\ \leq\ C\,2^{-\alpha k}\,\|v\|_{X^{2}},\quad k\in{\mathbb{N}}_{0}.$ (2.4)
@@
$\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{r;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{r}})^{2}<\infty$
and $(\sigma_{r;{\boldsymbol{s}}}^{-1})_{{\boldsymbol{s}}\in{\mathbb{F}}}\in\ell_{q_{r}}({\mathbb{F}})$ for some $0<q_{1}\leq q_{2}<\infty$. Define for $\xi>0$
@@
Hence, by the inequalities $q_{2}\alpha\leq 1$ and $\sigma_{2;{\boldsymbol{s}}}>1$ we derive that
@@
For the dimension of the space ${\mathcal{V}}(G(\xi))$ we have that
@@
The norms in the right hand side can be estimated using Parseval’s identity and the hypothesis of the theorem. Thus, for the norm $\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}$ we have that
$\begin{split}\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}^{2}\ &=\ \sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}\|v_{\boldsymbol{s}}\|_{X^{1}}^{2}\ =\ \sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\sigma_{1;{\boldsymbol{s}}}^{-2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\leq C\xi^{-2/q_{1}}.\end{split}$ (2.16)
@@ 3.1 Auxiliary results
Proof. Similarly to the proof of Lemma 3.3, it is sufficient to prove the lemma for $p=2$. Since in the case $\alpha\leq 1/q_{2}-1/2$, the formulas (2.7) and (3.5) define the same set $G(\xi)$ for $\xi>0$, from (2.12) follows the lemma for this case. Let us consider the case $\alpha>1/q_{2}-1/2$. Putting
@@
$\begin{split}\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}\|v_{\boldsymbol{s}}\|_{X^{1}}\,\|H_{\boldsymbol{s}}\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ =\ \sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}\|v_{\boldsymbol{s}}\|_{X^{1}}\\[6.45831pt] \ &\leq\ \left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}\sigma_{1;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\,\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\sigma_{1;{\boldsymbol{s}}}^{-(2-q_{1})}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\xi^{-(1/q_{1}-1/2)}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/2}\leq C\xi^{-(1/q_{1}-1/2)}.\end{split}$ (3.9)
@@
$\begin{split}\|v_{\xi}-{\mathcal{S}}_{G(\xi)}v\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\Big{\|}v_{\boldsymbol{s}}-\sum_{2^{k}\leq\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau}}\delta_{k}(v_{\boldsymbol{s}})\Big{\|}_{X^{1}}\,\|H_{\boldsymbol{s}}\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\[6.45831pt] \ &=\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\Big{\|}v_{\boldsymbol{s}}-P_{N}(v_{\boldsymbol{s}})\Big{\|}_{X^{1}}\ \leq\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}N^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\\[6.45831pt] \ &\leq\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}(\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau})^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\ \leq\ C\,\xi^{-\vartheta\alpha/\tau}\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{\alpha/\tau}\|v_{\boldsymbol{s}}\|_{X^{2}}\\[6.45831pt] \ &\leq\ C\,\xi^{-\vartheta\alpha/\tau}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}(\sigma_{2;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{2}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{2(\alpha/\tau-1)}\right)^{1/2}\\[6.45831pt] \ &=\ C\,\xi^{-(1/q_{1}-1/2)}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/2}\ \leq\ C\,\xi^{-(1/q_{1}-1/2)}\end{split}$
@@ 3.2 Interpolation approximation
be the Lebesgue constant, see, e.g., [26, Page 78]. We want to choose a sequence $(Y_{m})_{m=0}^{\infty}$ so that for some positive numbers $\tau$ and $C$, there holds the inequality
$\lambda_{m}(Y_{m})\ \leq\ (Cm+1)^{\kappa},\quad\forall m\in{\mathbb{N}}_{0}.$ (3.15)
@@
Assume that $(Y_{m})_{m=0}^{\infty}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$. Then for every $\varepsilon>0$, there exists a positive constant $C_{\varepsilon}$ independent of $m$ such that for every function $v$ on $\mathbb{R}$},
$\|\Delta^{{\rm I}}_{m}(v){\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})}\ \leq\ (C_{\varepsilon}m+1)^{\kappa+\varepsilon}\|v{\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})},\quad\forall m\in{\mathbb{N}}_{0},$ (3.17)
@@
$\|\Delta^{{\rm I}}_{m}(v){\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})}\ \leq\ 2(Cm+1)^{\kappa}\|v{\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})},\quad\forall m\in{\mathbb{N}}_{0},$
@@
Let $0<p\leq 2$. Let Assumption II hold for Hilbert spaces $X^{1}$ and $X^{2}$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$. Let $v\in{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})$ be represented by the series (2.5). Assume that for $r=1,2$ there exist increasing sequences $(\sigma_{r;{\boldsymbol{s}}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ of numbers strictly larger than $1$ such that
@@
$\theta:=\kappa+\varepsilon+5/4,\quad\lambda:=\max\left(C_{\infty,2},C_{2,\infty},C_{\varepsilon},1\right),$ (3.19)
$C_{\infty,2}$, $C_{2,\infty}$ are as in Lemma 3.1, $\varepsilon$ is arbitrary positive number and $C_{\varepsilon}$ is as in Lemma 3.7. For $\xi>0$, let $G(\xi)$ be the set defined as in (3.5).
Then there exists a constant $C$ such that for each $\xi>0$ , for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that for the interpolation operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})\to{\mathcal{V}}(G(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
$\beta:=\left(\frac{1}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{1}{q_{1}}-\frac{1}{q_{2}}.$ (3.21)
@@
Observe that $\Lambda_{k}(\xi)=\emptyset$ for all $k>k(\xi):=\lfloor\log_{2}\xi\rfloor$, and consequently, we have that
@@
Let us first consider the case $\alpha\leq 1/q_{2}-1/2$. From (3.24) it follows that
@@
Lemma 3.5 gives
Let us estimate the sum in the right-hand side of (3.25). We have that
@@
$\begin{split}\|\Delta^{{\rm I}}_{{\boldsymbol{s}}^{\prime}}(H_{\boldsymbol{s}})\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ &=\ \prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\|_{L_{2}({\mathbb{R}},\gamma)}\ =\ \prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{2}({\mathbb{R}})}\\[6.45831pt] \ &\leq\ p_{{\boldsymbol{s}}^{\prime}}(\theta,\lambda)\prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{\infty}({\mathbb{R}})}\end{split}$
@@
$\begin{split}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{\infty}({\mathbb{R}})}\ &\leq\ (1+C_{\varepsilon}s^{\prime}_{j})^{\kappa+\varepsilon}\,\|H_{s_{j}}\sqrt{g}\|_{L_{\infty}({\mathbb{R}})}\\ \ &\leq\ (1+C_{\varepsilon}s^{\prime}_{j})^{\kappa+\varepsilon}(1+C_{\infty,2}s_{j})^{1/4}\,\|H_{s_{j}}\sqrt{g}\|_{L_{2}({\mathbb{R}})}\\ \ &=\ (1+C_{\varepsilon}s^{\prime}_{j})^{\kappa+\varepsilon}(1+C_{\infty,2}s_{j})^{1/4}\end{split}$
@@
$\theta:=\kappa+\varepsilon+1/4,\quad\lambda:=\max\left(C_{\infty,2},C_{2,\infty},C_{\varepsilon}\right).$ (3.29)
@@
$\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ \leq\ \sum_{{\boldsymbol{s}}^{\prime}\in\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,p_{\boldsymbol{s}}(\theta,\lambda)\ \leq\ |R_{\boldsymbol{s}}|\,p_{\boldsymbol{s}}(\theta,\lambda)\ \leq\ p_{\boldsymbol{s}}(1,1)\,p_{\boldsymbol{s}}(\theta,\lambda).$
where $\theta$ and $\lambda$ are as in (3.29). Hence,
where $\theta$ and $\lambda$ are as in (3.19). Denote by $A(\xi)$ the sum in the right-hand side of (3.24), By using (3.30) and (2.4) we estimate the norm $\|A(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}$ as
$\begin{split}\|A(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ C\sum_{(k,{\boldsymbol{s}})\not\in G(\xi)}2^{-\alpha k}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ =\ C\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ \sum_{2^{k}>\xi\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}}2^{-\alpha k}\\[6.45831pt] \ &\leq\ C\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ (\xi\sigma_{2;{\boldsymbol{s}}}^{-q_{2}})^{-\alpha}\ \leq\ C\xi^{-\alpha}\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\sigma_{2;{\boldsymbol{s}}}^{q_{2}\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}.\end{split}$

### v14

# Sparse-grid polynomial interpolation approximation and integration for parametric and stochastic elliptic PDEs with lognormal inputs

@@ 1 Introduction
$\|P_{n}(w)\|_{V}\leq C\|w\|_{V},\quad\|w-P_{n}(w)\|_{V}\leq Cn^{-\alpha}\|w\|_{W},\quad\forall n\in{\mathbb{N}}_{0},\quad\forall w\in W.$ (1.5)
@@ 2 Linear approximation by truncated Hermite GPC expansion
$\|P_{n}(w)\|_{X^{1}}\leq C\|w\|_{X^{1}},\quad\|w-P_{n}(w)\|_{X^{1}}\leq Cn^{-\alpha}\|w\|_{X^{2}},\quad\forall n\in{\mathbb{N}}_{0},\quad\forall w\in X^{2}.$ (2.3)
@@
$\delta_{k}(w):=\ P_{2^{k}}(w)-P_{2^{k-1}}(w),\ k\in{\mathbb{N}},\quad\delta_{0}(w)=P_{0}(w).$
Under Assumption II, we can represent every $w\in X^{2}$ by the series
@@
converging in $X^{1}$ and satisfying the estimates
$\|\delta_{k}(w)\|_{X^{1}}\ \leq\ C\,\,\|w\|_{X^{1}},\quad\|\delta_{k}(w)\|_{X^{1}}\ \leq\ C\,2^{-\alpha k}\,\|w\|_{X^{2}},\quad k\in{\mathbb{N}}_{0}.$ (2.4)
@@
$\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{r;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{r}})^{2}<\infty$ (2.8)
and $(\sigma_{r;{\boldsymbol{s}}}^{-1})_{{\boldsymbol{s}}\in{\mathbb{F}}}\in\ell_{q_{r}}({\mathbb{F}})$ for some $0<q_{1}\leq q_{2}<\infty$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
Hence, by the inequalities $q_{2}\alpha\leq 1$ and $\sigma_{2;{\boldsymbol{s}}}>1$, and (2.8) we derive that
@@
For the dimension of the space ${\mathcal{V}}(G(\xi))$ we have by (2.8) that
@@
The norms in the right hand side can be estimated using Parseval’s identity and the hypothesis of the theorem. Thus, for the norm $\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}$ we have by (2.8) that
$\begin{split}\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}^{2}\ &=\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}^{2}\ =\ \sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\sigma_{1;{\boldsymbol{s}}}^{-2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\leq C\xi^{-2/q_{1}}.\end{split}$ (2.17)
@@ 3.1 Auxiliary results
Proof. Similarly to the proof of Lemma 3.3, it is sufficient to prove the lemma for $p=2$. Since in the case $\alpha\leq 1/q_{2}-1/2$, the formulas (2.7) and (3.5) define the same set $G(\xi)$ for $\xi>1$, from (2.13) follows the lemma for this case. Let us consider the case $\alpha>1/q_{2}-1/2$. Putting
@@
$\begin{split}\|v-v_{\xi}\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}>(\xi)^{1/q_{1}}}\|v_{\boldsymbol{s}}\|_{X^{1}}\,\|H_{\boldsymbol{s}}\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ =\ \sum_{\sigma_{1;{\boldsymbol{s}}}>(\xi)^{1/q_{1}}}\|v_{\boldsymbol{s}}\|_{X^{1}}\\[6.45831pt] \ &\leq\ \left(\sum_{\sigma_{1;{\boldsymbol{s}}}>(\xi)^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>(\xi)^{1/q_{1}}}\sigma_{1;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\,\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>(\xi)^{1/q_{1}}}\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\sigma_{1;{\boldsymbol{s}}}^{-(2-q_{1})}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\xi^{-(1/q_{1}-1/2)}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/2}\leq C(\xi)^{-(1/q_{1}-1/2)}.\end{split}$ (3.9)
@@
$\begin{split}\|v_{\xi}-{\mathcal{S}}_{G(\xi)}v\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\Big\|v_{\boldsymbol{s}}-\sum_{2^{k}\leq\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau}}\delta_{k}(v_{\boldsymbol{s}})\Big\|_{X^{1}}\,\|H_{\boldsymbol{s}}\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\[6.45831pt] \ &=\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\Big\|v_{\boldsymbol{s}}-P_{N}(v_{\boldsymbol{s}})\Big\|_{X^{1}}\ \leq\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}N^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\\[6.45831pt] \ &\leq\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}(\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau})^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\ \leq\ C\,\xi^{-\vartheta\alpha/\tau}\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{\alpha/\tau}\|v_{\boldsymbol{s}}\|_{X^{2}}\\[6.45831pt] \ &\leq\ C\,\xi^{-\vartheta\alpha/\tau}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}(\sigma_{2;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{2}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{2(\alpha/\tau-1)}\right)^{1/2}\\[6.45831pt] \ &=\ C\,\xi^{-(1/q_{1}-1/2)}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/2}\ \leq\ C\,\xi^{-(1/q_{1}-1/2)}.\end{split}$
@@ 3.2 Interpolation approximation
be the Lebesgue constant, see, e.g., [26, Page 78]. We want to choose a sequence $(Y_{m})_{m=0}^{\infty}$ so that for some positive numbers $\varkappa$ and $C$, there holds the inequality
$\lambda_{m}(Y_{m})\ \leq\ (Cm+1)^{\varkappa},\quad\forall m\in{\mathbb{N}}_{0}.$ (3.15)
@@
Assume that $(Y_{m})_{m=0}^{\infty}$ is a sequence satisfying the condition (3.15) for some positive numbers $\varkappa$ and $C$. Then for every $\varepsilon>0$, there exists a positive constant $C_{\varepsilon}$ independent of $m$ such that for every function $v$ on $\mathbb{R}$},
$\|\Delta^{{\rm I}}_{m}(v){\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})}\ \leq\ (C_{\varepsilon}m+1)^{\varkappa+\varepsilon}\|v{\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})},\quad\forall m\in{\mathbb{N}}_{0},$ (3.17)
@@
$\|\Delta^{{\rm I}}_{m}(v){\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})}\ \leq\ 2(Cm+1)^{\varkappa}\|v{\sqrt{g}}\|_{L_{\infty}({\mathbb{R}})},\quad\forall m\in{\mathbb{N}}_{0},$
@@
Let $0<p\leq 2$. Let Assumption II hold for Hilbert spaces $X^{1}$ and $X^{2}$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\varkappa$ and $C$. Let $v\in{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})$ be represented by the series (2.5). Assume that for $r=1,2$ there exist increasing sequences $(\sigma_{r;{\boldsymbol{s}}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ of numbers strictly larger than $1$ such that
@@
$\theta:=\varkappa+\varepsilon+5/4,\quad\lambda:=\max\left(C_{\infty,2},C_{2,\infty},C_{\varepsilon},1\right),$ (3.19)
$C_{\infty,2}$, $C_{2,\infty}$ are as in Lemma 3.1, $\varepsilon$ is arbitrary positive number and $C_{\varepsilon}$ is as in Lemma 3.7. For $\xi>1$, let $G(\xi)$ be the set defined as in (3.5).
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that for the interpolation operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})\to{\mathcal{V}}(G(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
$\beta:=\left(\frac{1}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\ \ \ \delta:=\frac{1}{q_{1}}-\frac{1}{q_{2}}.$ (3.21)
@@
For $\xi>1$, we put $k(\xi):=\lfloor\log_{2}\xi\rfloor$ if $\alpha\leq 1/q_{2}-1/2$, and $k(\xi):=\lfloor\vartheta\tau^{-1}\log_{2}\xi\rfloor$ if $\alpha>1/q_{2}-1/2$. Observe that $\Lambda_{k}(\xi)=\emptyset$ for all $k>k(\xi)$, and consequently, we have that
@@
Observe that for $k\leq k(\xi)$, if ${\boldsymbol{s}}\not\in\Lambda_{k}(\xi)$, then $(k,{\boldsymbol{s}})\not\in G(\xi)$. Hence, by (3.24) it follows that
@@
We need the following auxiliary bound for $\big\|I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}$,
where $\theta$ and $\lambda$ are as in (3.19). We have that
@@
$\begin{split}\|\Delta^{{\rm I}}_{{\boldsymbol{s}}^{\prime}}(H_{\boldsymbol{s}})\big\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ &=\ \prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\|_{L_{2}({\mathbb{R}},\gamma)}\ =\ \prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{2}({\mathbb{R}})}\\[6.45831pt] \ &\leq\ p_{{\boldsymbol{s}}^{\prime}}(\theta,\lambda)\prod_{j=1}^{J}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{\infty}({\mathbb{R}})},\end{split}$
@@
$\begin{split}\|\Delta^{{\rm I}}_{s^{\prime}_{j}}(H_{s_{j}})\sqrt{g}\|_{L_{\infty}({\mathbb{R}})}\ &\leq\ (1+C_{\varepsilon}s^{\prime}_{j})^{\varkappa+\varepsilon}\,\|H_{s_{j}}\sqrt{g}\|_{L_{\infty}({\mathbb{R}})}\\ \ &\leq\ (1+C_{\varepsilon}s^{\prime}_{j})^{\varkappa+\varepsilon}(1+C_{\infty,2}s_{j})^{1/4}\,\|H_{s_{j}}\sqrt{g}\|_{L_{2}({\mathbb{R}})}\\ \ &=\ (1+C_{\varepsilon}s^{\prime}_{j})^{\varkappa+\varepsilon}(1+C_{\infty,2}s_{j})^{1/4}\end{split}$
@@
$\theta:=\varkappa+\varepsilon+1/4,\quad\lambda:=\max\left(C_{\infty,2},C_{2,\infty},C_{\varepsilon}\right).$ (3.29)
@@
$\big\|I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ \leq\ \sum_{{\boldsymbol{s}}^{\prime}\in\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,p_{\boldsymbol{s}}(\theta,\lambda)\ \leq\ \min\big(|\Lambda_{k}(\xi)|,|R_{\boldsymbol{s}}|\big)\,p_{\boldsymbol{s}}(\theta,\lambda),$ (3.30)
where $\theta$ and $\lambda$ are as in (3.29). Hence, by the inequality $|R_{\boldsymbol{s}}|\leq p_{\boldsymbol{s}}(1,1)$ we get (3.26).
From (3.25) and (3.26) it follows that
$\big\|v-{\mathcal{I}}_{G(\xi)}v\big\|_{{\mathcal{L}}_{2}(X^{1})}\ \leq\ \big\|v-{\mathcal{S}}_{G(\xi)}v\big\|_{{\mathcal{L}}_{2}(X^{1})}+A(\xi).$ (3.31)
where
$A(\xi):=\ \sum_{(k,{\boldsymbol{s}})\not\in G(\xi)}\|\delta_{k}(v_{\boldsymbol{s}})\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda),$ (3.32)
where $\theta$ and $\lambda$ are as in (3.19). In the next step, we employ the inequality (3.31) to establish bounds for $\big\|v-{\mathcal{I}}_{G(\xi)}v\big\|_{{\mathcal{L}}_{2}(X^{1})}$.
Let us first consider the case $\alpha\leq 1/q_{2}-1/2$. Lemma 3.5 gives
Let us estimate the term $A(\xi)$ in the right-hand side of (3.31). By using (2.4) we derive that
$\begin{split}A(\xi)\ &\leq\ C\sum_{(k,{\boldsymbol{s}})\not\in G(\xi)}2^{-\alpha k}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ =\ C\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ \sum_{2^{k}>\xi\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}}2^{-\alpha k}\\[6.45831pt] \ &\leq\ C\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\|v_{\boldsymbol{s}}\|_{X^{2}}\ (\xi\sigma_{2;{\boldsymbol{s}}}^{-q_{2}})^{-\alpha}\ \leq\ C\xi^{-\alpha}\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(\theta,\lambda)\sigma_{2;{\boldsymbol{s}}}^{q_{2}\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}.\end{split}$